OpenStudy (anonymous):

If f(x) = 6 – 5x and g(x) = 4x – 1, evaluate f(x) – g(x) for x = –2.

OpenStudy (anonymous):

@iGreen

OpenStudy (12man):

f(x) = 6 – 5x g(x) = 4x – 1 f(x) – g(x) for x = –2 (6 - 5(-2)) - (4(-2) - 1) = 16 - (-9) = 16 + 9 = 25 (D)

OpenStudy (anonymous):

just plug it in... 6 - 5(-2) = 16

OpenStudy (anonymous):

6 -(-10) 6 + 10 16

OpenStudy (anonymous):

it is D

OpenStudy (anonymous):

25

OpenStudy (anonymous):

do the same for g(x)

OpenStudy (12man):

yes

OpenStudy (anonymous):

4(-2) - 1 = -9

OpenStudy (anonymous):

16 -(-9) 16 + 9 25

OpenStudy (12man):

yes 25

OpenStudy (anonymous):

ya it is 25

OpenStudy (ageta):

i think it is 25

OpenStudy (ageta):

right